OpenStudy (anonymous):

for the given equation list the intercepts and test for symmetry y=-8x/x^2+4

OpenStudy (campbell_st):

the only intercept is when -8x = 0 which is when x = 0 the the x intercept is 0, and y- intercept is 0. the function is odd since \[f(-x) = \frac{-8(-x)}{(-x)^2 + 4} = \frac{8x}{x^2 + 4} = -f(x)\] so the curve has rotational symmetry about (0, 0)

OpenStudy (anonymous):

Thank you!!!!